# MiniLLM: Knowledge Distillation of Large Language Models

### v5

@@ C.1 GPT-J as the Teacher Model
We present the evaluation results when using GPT-J as the teacher model and GPT-2-760M, GPT-2-1.5B, and GPT-Neo-2.7B [gpt-neo] as the student models in Table 6. MiniLLM outperforms the baselines in most cases.
@@ C.4 More Ablation Studies
In Table 16, we study the effect of adding the pre-training loss in Algorithm 1 by comparing MiniLLM with its variant where the language modeling loss on the pre-training corpus is removed (w/o $\mathcal{L}_{\text{PT}}$). We have a similar observation as [instruct-gpt] that adding the pre-training loss helps to preserve the abilities on canonical NLP tasks while keeping the performance on instruction-following tasks nearly unchanged.
@@ Appendix D Cases
MiniLLM
w/o teacher-mixed
We to know
@@
MiniLLM
w/o teacher-mixed
SAMSUNGC 980 Pro SSD 2TB PCIe NVME Gen 4 Gaming M.II Internal Solid State Hard Drive memory Card, Max Speed, Thermal Control, AP1500K

### v6

@@ C.1 GPT-J as the Teacher Model
We present the evaluation results when using GPT-J as the teacher model and GPT-2-760M, GPT-2-1.5B, and GPT-Neo-2.7B [7] as the student models in Table 6. MiniLLM outperforms the baselines in most cases.
@@ C.4 More Ablation Studies
In Table 16, we study the effect of adding the pre-training loss in Algorithm 1 by comparing MiniLLM with its variant where the language modeling loss on the pre-training corpus is removed (w/o $\mathcal{L}_{\text{PT}}$). We have a similar observation as [51] that adding the pre-training loss helps to preserve the abilities on canonical NLP tasks while keeping the performance on instruction-following tasks nearly unchanged.
@@ Appendix D Cases
MiniLLM w/o teacher-mixed We to know
@@
MiniLLM w/o teacher-mixed SAMSUNGC 980 Pro SSD 2TB PCIe NVME Gen 4 Gaming M.II Internal Solid State Hard Drive memory Card, Max Speed, Thermal Control, AP1500K